Circular No. 7619

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bb IN IC 4319 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001bb, obtained by P. Berlind on May 2.27 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ic supernova near maximum light. The spectrum is very similar to that of SN 1994I at 4 days before maximum (see Figure 1 of Filippenko et al. 1995, Ap.J. 450, L11), although SN 2001bb appears to be more reddened. It does not appear to be similar to the spectra of hypernovae (cf. IAUC 7616) such as SN 1998bw (Iwamoto et al. 1998, Nature 395, 672) or SN 1997ef (Iwamoto et al. 2000, Ap.J. 534, 660).

GU SAGITTARII E. O. Waagen, American Association of Variable Star Observers, re-

ports that, according to the AAVSO International Database, the last fading of this R Crb-type variable began between Dec. 1987 and Feb. 1988 (when it reached minimum visual mag ∼ 16.3), before it returned to full maximum at m_{v} = 10.4 in Feb. 1993. Since then, GU Sgr has been observed around maximum until it was found to have faded this past month. Recent magnitude estimates reported to the AAVSO (visual unless otherwise noted): Apr. 15.27 UT, [12.4 (S. Otero, Buenos Aires, Argentina); 16.330, 14.3 (S. O’Connor, Montreal North, QC; CCD V); 20.356, 14.3 (O’Connor); 25.622, [14.2 (R. Stubbings, Druin, Victoria, Australia); 30.847, [14.5 (A. Pearce, Nedlands, W. Australia); May 2.835, [14.6 (Pearce); 3.848, [14.5 (Pearce).

2000 GN_{171}

S. S. Sheppard, University of Hawaii (UH), writes: “R-band CCD photometry with the UH 2.2-m telescope of the transneptunian object 2000 GN_{171} (ephemeris on MPEC 2001-G04) during Apr. 20–25 shows a singlepeaked rotational lightcurve. It has a peak-to-peak variation of 0.6 magnitude and a period of 4.15 ± 0.05 hr, though an alias period of ∼ 3.5 hr is possible. The object could be elongated with a double-peak of ∼ 8.3 hr.”

COMET 45P/HONDA-MRKOS-PAJDUŠÁKOVÁ Visual m_{1} estimates: Apr. 12.12 UT, 9.4 (A. Hale, Cloudcroft, NM,

0.2-m reflector); 13.84, 9.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 22.47, 10.2 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); 26.84, 10.1 (W. Hasubick, Buchloe, Germany, 0.44-m reflector).

2001 May 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green